# Research Paper

### v1

@@ A.1 PyCCF
where and represent the sets of photometric and spectroscopic amplitudes respectively, and angled brackets indicate an inner product, approximated by a sum.777, where the overline represents the mean. The physical lag is then taken to be the delay that maximises this correlation, while the uncertainties are estimated by ‘bootstrapping’: recovering the peak correlation-lag from many realisations generated by randomly sub-sampling the observations and then re-sampling these observations within their measurement uncertainties.
@@ A.2 JAVELIN
Equation 14 requires the inversion of a large matrix of rank . Though the near-diagonal shape of equation 13 ameliorates this cost somewhat, this is still a computationally expensive process. To reduce the number of evaluations, JAVELIN uses the Monte Carlo Markov Chain (MCMC) approach using the python-based package emcee (emcee-Foreman_Mackey_2013), which uses the Affine-Invariant Ensemble Sampler (AffineInvariant-Goodman_2010) to direct the bulk of its samples to high likelihood regions of parameter space. The seasonal windowing function of equatorial surveys, and the resulting aliasing problem (see Section 2.2 for details) can give rise to multi-modal distributions along the lag axis of parameter-space (e.g. CIV_Grier_2019), which can result in poor efficiency and inaccurate posterior distributions from emcee. Gaussian-process-based modelling is also sensitive to outlier data-points or under-estimated measurement uncertainties, which can spuriously over-constrain the light curve and disrupt the lag recovery results.
@@ Appendix B OzDES Reverberation Mapping Results
| Source ID | Redshift |
| 0.127 | ||||||||||||||||||||
| 0.237 | ||||||||||||||||||||
| 0.310 | ||||||||||||||||||||
| 0.314 | ||||||||||||||||||||
| 0.332 | ||||||||||||||||||||
| 0.354 | ||||||||||||||||||||
| 0.644 | ||||||||||||||||||||
@@
| Source ID | Redshift |
| 0.840 | ||||||||||||||||||||
| 0.890 | ||||||||||||||||||||
| 1.000 | ||||||||||||||||||||
| 1.070 | ||||||||||||||||||||
| 1.150 | ||||||||||||||||||||
| 1.240 | ||||||||||||||||||||
| 1.370 | ||||||||||||||||||||
| 1.420 | ||||||||||||||||||||
| 1.430 | ||||||||||||||||||||
| 1.440 | ||||||||||||||||||||
| 1.480 | ||||||||||||||||||||
| 1.490 | ||||||||||||||||||||
| 1.490 | ||||||||||||||||||||
| 1.530 | ||||||||||||||||||||
| 1.530 | ||||||||||||||||||||
| 1.620 | ||||||||||||||||||||
| 1.640 | ||||||||||||||||||||
| 1.640 | ||||||||||||||||||||
| 1.640 | ||||||||||||||||||||
| 1.650 | ||||||||||||||||||||
| 1.660 | ||||||||||||||||||||
| 1.680 | ||||||||||||||||||||
| 1.750 | ||||||||||||||||||||
| 1.790 | ||||||||||||||||||||
@@
| Source ID | Redshift |
| 1.922 | ||||||||||||||||||||
| 1.928 | ||||||||||||||||||||
| 1.936 | ||||||||||||||||||||
| 1.952 | ||||||||||||||||||||
| 1.988 | ||||||||||||||||||||
| 2.031 | ||||||||||||||||||||
| 2.041 | ||||||||||||||||||||
| 2.092 | ||||||||||||||||||||
| 2.096 | ||||||||||||||||||||
| 2.150 | ||||||||||||||||||||
| 2.188 | ||||||||||||||||||||
| 2.219 | ||||||||||||||||||||
| 2.257 | ||||||||||||||||||||
| 2.283 | ||||||||||||||||||||
| 2.315 | ||||||||||||||||||||
| 2.328 | ||||||||||||||||||||
| 2.384 | ||||||||||||||||||||
| 2.425 | ||||||||||||||||||||
| 2.452 | ||||||||||||||||||||
| 2.459 | - | - | - | |||||||||||||||||
| 2.462 | ||||||||||||||||||||
| 2.492 | ||||||||||||||||||||
| 2.500 | ||||||||||||||||||||
| 2.563 | ||||||||||||||||||||
| 2.593 | ||||||||||||||||||||
| 2.677 | ||||||||||||||||||||
| 2.745 | ||||||||||||||||||||
| 2.773 | - | - | - | |||||||||||||||||
@@ Appendix C OzDES Single Epoch Results
| Source ID | Redshift | Line |
| 0.040 | H | |||||||||||||||||||||||||
| 0.180 | H | |||||||||||||||||||||||||
| 0.220 | H | |||||||||||||||||||||||||
| 0.250 | H | |||||||||||||||||||||||||
| 0.310 | H | |||||||||||||||||||||||||
| 0.430 | H | |||||||||||||||||||||||||
| 0.490 | H | |||||||||||||||||||||||||
| 0.510 | H | |||||||||||||||||||||||||
| 0.540 | H | |||||||||||||||||||||||||
| 0.550 | H | |||||||||||||||||||||||||
| 0.570 | H | |||||||||||||||||||||||||
| 0.590 | H | |||||||||||||||||||||||||
| 0.610 | H | |||||||||||||||||||||||||
| 0.620 | H | |||||||||||||||||||||||||
| 0.640 | H | |||||||||||||||||||||||||
| 0.650 | H | |||||||||||||||||||||||||
| 0.660 | H | |||||||||||||||||||||||||
| 0.670 | H | |||||||||||||||||||||||||
| 0.670 | H | |||||||||||||||||||||||||
| 0.710 | MgII | |||||||||||||||||||||||||
| 0.720 | MgII | |||||||||||||||||||||||||
| 0.720 | MgII | |||||||||||||||||||||||||
| 0.730 | MgII | |||||||||||||||||||||||||
| 0.730 | MgII | |||||||||||||||||||||||||
| 0.730 | MgII | |||||||||||||||||||||||||
| 0.760 | MgII | |||||||||||||||||||||||||
| 0.770 | MgII | |||||||||||||||||||||||||
| 0.790 | MgII | |||||||||||||||||||||||||
| 0.790 | MgII | |||||||||||||||||||||||||
| 0.800 | MgII | |||||||||||||||||||||||||
| 0.810 | MgII | |||||||||||||||||||||||||
| 0.820 | MgII | |||||||||||||||||||||||||
| 0.830 | MgII | |||||||||||||||||||||||||
| 0.840 | MgII | |||||||||||||||||||||||||
| 0.860 | MgII | |||||||||||||||||||||||||
| 0.860 | MgII | |||||||||||||||||||||||||
| 0.870 | MgII | |||||||||||||||||||||||||
| 0.880 | MgII | |||||||||||||||||||||||||
| 0.890 | MgII | |||||||||||||||||||||||||
| 0.900 | MgII | |||||||||||||||||||||||||
| 0.900 | MgII | |||||||||||||||||||||||||
| 0.910 | MgII | |||||||||||||||||||||||||
| 0.910 | MgII | |||||||||||||||||||||||||
| 0.920 | MgII | |||||||||||||||||||||||||
| 0.920 | MgII | |||||||||||||||||||||||||
| 0.930 | MgII | |||||||||||||||||||||||||
| 0.930 | MgII | |||||||||||||||||||||||||
| 0.940 | MgII | |||||||||||||||||||||||||
| 0.940 | MgII | |||||||||||||||||||||||||
| 0.940 | MgII | |||||||||||||||||||||||||
| 0.940 | MgII | |||||||||||||||||||||||||
| 0.950 | MgII | |||||||||||||||||||||||||
| 0.950 | MgII | |||||||||||||||||||||||||
| 0.950 | MgII | |||||||||||||||||||||||||
| 0.960 | MgII | |||||||||||||||||||||||||
| 0.960 | MgII | |||||||||||||||||||||||||
| 0.960 | MgII | |||||||||||||||||||||||||
| 0.960 | MgII | |||||||||||||||||||||||||
| 0.960 | MgII | |||||||||||||||||||||||||
| 0.960 | MgII | |||||||||||||||||||||||||
| 0.970 | MgII | |||||||||||||||||||||||||
| 0.970 | MgII | |||||||||||||||||||||||||
| 0.970 | MgII | |||||||||||||||||||||||||
| 0.970 | MgII | |||||||||||||||||||||||||
| 0.980 | MgII | |||||||||||||||||||||||||
| 0.980 | MgII | |||||||||||||||||||||||||
| 0.980 | MgII | |||||||||||||||||||||||||
| 0.980 | MgII | |||||||||||||||||||||||||
| 0.980 | MgII | |||||||||||||||||||||||||
| 0.990 | MgII | |||||||||||||||||||||||||
| 0.990 | MgII | |||||||||||||||||||||||||
| 1.010 | MgII | |||||||||||||||||||||||||
| 1.010 | MgII | |||||||||||||||||||||||||
| 1.020 | MgII | |||||||||||||||||||||||||
| 1.050 | MgII | |||||||||||||||||||||||||
| 1.070 | MgII | |||||||||||||||||||||||||
| 1.080 | MgII | |||||||||||||||||||||||||
| 1.090 | MgII | |||||||||||||||||||||||||
| 1.090 | MgII | |||||||||||||||||||||||||
| 1.090 | MgII | |||||||||||||||||||||||||
| 1.090 | MgII | |||||||||||||||||||||||||
| 1.090 | MgII | |||||||||||||||||||||||||
| 1.090 | MgII | |||||||||||||||||||||||||
| 1.110 | MgII | |||||||||||||||||||||||||
| 1.110 | MgII | |||||||||||||||||||||||||
| 1.120 | MgII | |||||||||||||||||||||||||
| 1.130 | MgII | |||||||||||||||||||||||||
| 1.140 | MgII | |||||||||||||||||||||||||
| 1.160 | MgII | |||||||||||||||||||||||||
| 1.160 | MgII | |||||||||||||||||||||||||
| 1.160 | MgII | |||||||||||||||||||||||||
| 1.170 | MgII | |||||||||||||||||||||||||
| 1.170 | MgII | |||||||||||||||||||||||||
| 1.180 | MgII | |||||||||||||||||||||||||
| 1.190 | MgII | |||||||||||||||||||||||||
| 1.190 | MgII | |||||||||||||||||||||||||
| 1.200 | MgII | |||||||||||||||||||||||||
| 1.200 | MgII | |||||||||||||||||||||||||
| 1.200 | MgII | |||||||||||||||||||||||||
| 1.210 | MgII | |||||||||||||||||||||||||
| 1.220 | MgII | |||||||||||||||||||||||||
| 1.220 | MgII | |||||||||||||||||||||||||
| 1.230 | MgII | |||||||||||||||||||||||||
| 1.230 | MgII | |||||||||||||||||||||||||
| 1.230 | MgII | |||||||||||||||||||||||||
| 1.240 | MgII | |||||||||||||||||||||||||
| 1.240 | MgII | |||||||||||||||||||||||||
| 1.250 | MgII | |||||||||||||||||||||||||
| 1.250 | MgII | |||||||||||||||||||||||||
| 1.250 | MgII | |||||||||||||||||||||||||
| 1.250 | MgII | |||||||||||||||||||||||||
| 1.260 | MgII | |||||||||||||||||||||||||
| 1.270 | MgII | |||||||||||||||||||||||||
| 1.280 | MgII | |||||||||||||||||||||||||
| 1.280 | MgII | |||||||||||||||||||||||||
| 1.290 | MgII | |||||||||||||||||||||||||
| 1.290 | MgII | |||||||||||||||||||||||||
| 1.300 | MgII | |||||||||||||||||||||||||
| 1.300 | MgII | |||||||||||||||||||||||||
| 1.320 | MgII | |||||||||||||||||||||||||
| 1.320 | MgII | |||||||||||||||||||||||||
| 1.330 | MgII | |||||||||||||||||||||||||
| 1.340 | MgII | |||||||||||||||||||||||||
| 1.340 | MgII | |||||||||||||||||||||||||
| 1.350 | MgII | |||||||||||||||||||||||||
| 1.350 | MgII | |||||||||||||||||||||||||
| 1.350 | MgII | |||||||||||||||||||||||||
| 1.360 | MgII | |||||||||||||||||||||||||
| 1.360 | MgII | |||||||||||||||||||||||||
| 1.370 | MgII | |||||||||||||||||||||||||
| 1.370 | MgII | |||||||||||||||||||||||||
| 1.370 | MgII | |||||||||||||||||||||||||
| 1.370 | MgII | |||||||||||||||||||||||||
| 1.380 | MgII | |||||||||||||||||||||||||
| 1.380 | MgII | |||||||||||||||||||||||||
| 1.390 | MgII | |||||||||||||||||||||||||
| 1.390 | MgII | |||||||||||||||||||||||||
| 1.390 | MgII | |||||||||||||||||||||||||
| 1.390 | MgII | |||||||||||||||||||||||||
| 1.390 | MgII | |||||||||||||||||||||||||
| 1.400 | MgII | |||||||||||||||||||||||||
| 1.400 | MgII | |||||||||||||||||||||||||
| 1.420 | MgII | |||||||||||||||||||||||||
| 1.420 | MgII | |||||||||||||||||||||||||
| 1.420 | MgII | |||||||||||||||||||||||||
| 1.420 | MgII | |||||||||||||||||||||||||
| 1.430 | MgII | |||||||||||||||||||||||||
| 1.430 | MgII | |||||||||||||||||||||||||
| 1.430 | MgII | |||||||||||||||||||||||||
| 1.440 | MgII | |||||||||||||||||||||||||
| 1.450 | MgII | |||||||||||||||||||||||||
| 1.450 | MgII | |||||||||||||||||||||||||
| 1.450 | MgII | |||||||||||||||||||||||||
| 1.460 | MgII | |||||||||||||||||||||||||
| 1.460 | MgII | |||||||||||||||||||||||||
| 1.460 | MgII | |||||||||||||||||||||||||
| 1.470 | MgII | |||||||||||||||||||||||||
| 1.470 | MgII | |||||||||||||||||||||||||
| 1.470 | MgII | |||||||||||||||||||||||||
| 1.470 | MgII | |||||||||||||||||||||||||
| 1.480 | MgII | |||||||||||||||||||||||||
| 1.490 | MgII | |||||||||||||||||||||||||
| 1.490 | MgII | |||||||||||||||||||||||||
| 1.490 | MgII | |||||||||||||||||||||||||
| 1.490 | MgII | |||||||||||||||||||||||||
| 1.490 | MgII | |||||||||||||||||||||||||
| 1.500 | MgII | |||||||||||||||||||||||||
| 1.500 | MgII | |||||||||||||||||||||||||
| 1.530 | MgII | |||||||||||||||||||||||||
| 1.530 | MgII | |||||||||||||||||||||||||
| 1.530 | MgII | |||||||||||||||||||||||||
| 1.540 | MgII | |||||||||||||||||||||||||
| 1.550 | MgII | |||||||||||||||||||||||||
| 1.550 | MgII | |||||||||||||||||||||||||
| 1.560 | MgII | |||||||||||||||||||||||||
| 1.560 | MgII | |||||||||||||||||||||||||
| 1.570 | MgII | |||||||||||||||||||||||||
| 1.580 | MgII | |||||||||||||||||||||||||
| 1.590 | MgII | |||||||||||||||||||||||||
| 1.590 | MgII | |||||||||||||||||||||||||
| 1.600 | MgII | |||||||||||||||||||||||||
| 1.600 | MgII | |||||||||||||||||||||||||
| 1.610 | MgII | |||||||||||||||||||||||||
| 1.610 | MgII | |||||||||||||||||||||||||
| 1.620 | MgII | |||||||||||||||||||||||||
| 1.620 | MgII | |||||||||||||||||||||||||
| 1.630 | MgII | |||||||||||||||||||||||||
| 1.640 | MgII | |||||||||||||||||||||||||
| 1.650 | MgII | |||||||||||||||||||||||||
| 1.650 | MgII | |||||||||||||||||||||||||
| 1.680 | MgII | |||||||||||||||||||||||||
| 1.680 | MgII | |||||||||||||||||||||||||
| 1.680 | MgII | |||||||||||||||||||||||||
| 1.680 | MgII | |||||||||||||||||||||||||
| 1.690 | MgII | |||||||||||||||||||||||||
| 1.690 | MgII | |||||||||||||||||||||||||
| 1.690 | MgII | |||||||||||||||||||||||||
| 1.700 | MgII | |||||||||||||||||||||||||
| 1.700 | MgII | |||||||||||||||||||||||||
| 1.710 | MgII | |||||||||||||||||||||||||
| 1.720 | MgII | |||||||||||||||||||||||||
| 1.730 | MgII | |||||||||||||||||||||||||
| 1.730 | MgII | |||||||||||||||||||||||||
| 1.750 | MgII | |||||||||||||||||||||||||
| 1.760 | MgII | |||||||||||||||||||||||||
| 1.800 | MgII | |||||||||||||||||||||||||
| 1.800 | CIV | |||||||||||||||||||||||||
| 1.830 | MgII | |||||||||||||||||||||||||
| 1.830 | CIV | |||||||||||||||||||||||||
| 1.830 | MgII | |||||||||||||||||||||||||
| 1.830 | CIV | |||||||||||||||||||||||||
| 1.830 | MgII | |||||||||||||||||||||||||
| 1.830 | CIV | |||||||||||||||||||||||||
| 1.850 | MgII | |||||||||||||||||||||||||
| 1.850 | CIV | |||||||||||||||||||||||||
| 1.860 | MgII | |||||||||||||||||||||||||
| 1.870 | MgII | |||||||||||||||||||||||||
| 1.870 | CIV | |||||||||||||||||||||||||
| 1.880 | MgII | |||||||||||||||||||||||||
| 1.890 | MgII | |||||||||||||||||||||||||
| 1.890 | CIV | |||||||||||||||||||||||||
| 1.900 | MgII | |||||||||||||||||||||||||
| 1.900 | CIV | |||||||||||||||||||||||||
| 1.910 | MgII | |||||||||||||||||||||||||
| 1.920 | CIV | |||||||||||||||||||||||||
| 1.920 | MgII | |||||||||||||||||||||||||
| 1.930 | CIV | |||||||||||||||||||||||||
| 1.970 | CIV | |||||||||||||||||||||||||
| 1.990 | CIV | |||||||||||||||||||||||||
| 2.030 | CIV | |||||||||||||||||||||||||
| 2.060 | CIV | |||||||||||||||||||||||||
| 2.070 | CIV | |||||||||||||||||||||||||
| 2.080 | CIV | |||||||||||||||||||||||||
| 2.100 | CIV | |||||||||||||||||||||||||
| 2.110 | CIV | |||||||||||||||||||||||||
| 2.200 | CIV | |||||||||||||||||||||||||
| 2.220 | CIV | |||||||||||||||||||||||||
| 2.230 | CIV | |||||||||||||||||||||||||
| 2.240 | CIV | |||||||||||||||||||||||||
| 2.250 | CIV | |||||||||||||||||||||||||
| 2.250 | CIV | |||||||||||||||||||||||||
| 2.260 | CIV | |||||||||||||||||||||||||
| 2.290 | CIV | |||||||||||||||||||||||||
| 2.390 | CIV | |||||||||||||||||||||||||
| 2.440 | CIV | |||||||||||||||||||||||||

### v2

@@ A.1 PyCCF
where and represent the sets of photometric and spectroscopic amplitudes respectively, and angled brackets indicate an inner product, approximated by a sum.888, where the overline represents the mean. The physical lag is then taken to be the delay that maximises this correlation, while the uncertainties are estimated by ‘bootstrapping’: recovering the peak correlation-lag from many realisations generated by randomly sub-sampling the observations and then re-sampling these observations within their measurement uncertainties.
@@ A.2 JAVELIN
Equation 14 requires the inversion of a large matrix of rank . Though the near-diagonal shape of equation 13 ameliorates this cost somewhat, this is still a computationally expensive process. To reduce the number of evaluations, JAVELIN uses the MCMC approach using the python-based package emcee (emcee-Foreman_Mackey_2013), which uses the Affine-Invariant Ensemble Sampler (AffineInvariant-Goodman_2010) to direct the bulk of its samples to high likelihood regions of parameter space. The seasonal windowing function of equatorial surveys, and the resulting aliasing problem (see Section 2.2 for details) can give rise to multi-modal distributions along the lag axis of parameter-space (e.g. CIV_Grier_2019), which can result in poor efficiency and inaccurate posterior distributions from emcee. Gaussian-process-based modelling is also sensitive to outlier data-points or under-estimated measurement uncertainties, which can spuriously over-constrain the light curve and disrupt the lag recovery results.
@@ Appendix B OzDES Reverberation Mapping Results
| Source ID | Redshift | Monochromatic Luminosity | Rest-Frame Lag (Days) | Velocity Dispersion (km/s) | Mass () | Dimensionless Accretion Rate |
| 0.127 | ||||||
| 0.237 | ||||||
| 0.310 | ||||||
| 0.314 | ||||||
| 0.332 | ||||||
| 0.354 | ||||||
| 0.644 | ||||||
@@
| Source ID | Redshift | Monochromatic Luminosity | Rest-Frame Lag (Days) | Velocity Dispersion (km/s) | Mass () | Dimensionless Accretion Rate |
| 0.840 | ||||||
| 0.890 | ||||||
| 1.000 | ||||||
| 1.070 | ||||||
| 1.150 | ||||||
| 1.240 | ||||||
| 1.370 | ||||||
| 1.420 | ||||||
| 1.430 | ||||||
| 1.440 | ||||||
| 1.480 | ||||||
| 1.490 | ||||||
| 1.490 | ||||||
| 1.530 | ||||||
| 1.530 | ||||||
| 1.620 | ||||||
| 1.640 | ||||||
| 1.640 | ||||||
| 1.640 | ||||||
| 1.650 | ||||||
| 1.660 | ||||||
| 1.680 | ||||||
| 1.750 | ||||||
| 1.790 | ||||||
@@
| Source ID | Redshift | Monochromatic Luminosity | Rest-Frame Lag (Days) | Velocity Dispersion (km/s) | Mass () | Dimensionless Accretion Rate |
| 1.922 | ||||||
| 1.928 | ||||||
| 1.936 | ||||||
| 1.952 | ||||||
| 1.988 | ||||||
| 2.031 | ||||||
| 2.041 | ||||||
| 2.092 | ||||||
| 2.096 | ||||||
| 2.150 | ||||||
| 2.188 | ||||||
| 2.219 | ||||||
| 2.257 | ||||||
| 2.283 | ||||||
| 2.315 | ||||||
| 2.328 | ||||||
| 2.384 | ||||||
| 2.425 | ||||||
| 2.452 | ||||||
| 2.459 | - | - | - | |||
| 2.462 | ||||||
| 2.492 | ||||||
| 2.500 | ||||||
| 2.563 | ||||||
| 2.593 | ||||||
| 2.677 | ||||||
| 2.745 | ||||||
| 2.773 | - | - | - | |||
@@ Appendix C OzDES Single Epoch Results
| Source ID | Redshift | Line | Monochromatic Luminosity | Single Epoch Lag (Days) | Velocity Dispersion (km/s) | Single Epoch Mass () | Dimensionless Accretion Rate |
| 0.040 | H | ||||||
| 0.180 | H | ||||||
| 0.220 | H | ||||||
| 0.250 | H | ||||||
| 0.310 | H | ||||||
| 0.430 | H | ||||||
| 0.490 | H | ||||||
| 0.510 | H | ||||||
| 0.540 | H | ||||||
| 0.550 | H | ||||||
| 0.570 | H | ||||||
| 0.590 | H | ||||||
| 0.610 | H | ||||||
| 0.620 | H | ||||||
| 0.640 | H | ||||||
| 0.650 | H | ||||||
| 0.660 | H | ||||||
| 0.670 | H | ||||||
| 0.670 | H | ||||||
| 0.710 | MgII | ||||||
| 0.720 | MgII | ||||||
| 0.720 | MgII | ||||||
| 0.730 | MgII | ||||||
| 0.730 | MgII | ||||||
| 0.730 | MgII | ||||||
| 0.760 | MgII | ||||||
| 0.770 | MgII | ||||||
| 0.790 | MgII | ||||||
| 0.790 | MgII | ||||||
| 0.800 | MgII | ||||||
| 0.810 | MgII | ||||||
| 0.820 | MgII | ||||||
| 0.830 | MgII | ||||||
| 0.840 | MgII | ||||||
| 0.860 | MgII | ||||||
| 0.860 | MgII | ||||||
| 0.870 | MgII | ||||||
| 0.880 | MgII | ||||||
| 0.890 | MgII | ||||||
| 0.900 | MgII | ||||||
| 0.900 | MgII | ||||||
| 0.910 | MgII | ||||||
| 0.910 | MgII | ||||||
| 0.920 | MgII | ||||||
| 0.920 | MgII | ||||||
| 0.930 | MgII | ||||||
| 0.930 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.940 | MgII | ||||||
| 0.950 | MgII | ||||||
| 0.950 | MgII | ||||||
| 0.950 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.960 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.970 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.980 | MgII | ||||||
| 0.990 | MgII | ||||||
| 0.990 | MgII | ||||||
| 1.010 | MgII | ||||||
| 1.010 | MgII | ||||||
| 1.020 | MgII | ||||||
| 1.050 | MgII | ||||||
| 1.070 | MgII | ||||||
| 1.080 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.090 | MgII | ||||||
| 1.110 | MgII | ||||||
| 1.110 | MgII | ||||||
| 1.120 | MgII | ||||||
| 1.130 | MgII | ||||||
| 1.140 | MgII | ||||||
| 1.160 | MgII | ||||||
| 1.160 | MgII | ||||||
| 1.160 | MgII | ||||||
| 1.170 | MgII | ||||||
| 1.170 | MgII | ||||||
| 1.180 | MgII | ||||||
| 1.190 | MgII | ||||||
| 1.190 | MgII | ||||||
| 1.200 | MgII | ||||||
| 1.200 | MgII | ||||||
| 1.200 | MgII | ||||||
| 1.210 | MgII | ||||||
| 1.220 | MgII | ||||||
| 1.220 | MgII | ||||||
| 1.230 | MgII | ||||||
| 1.230 | MgII | ||||||
| 1.230 | MgII | ||||||
| 1.240 | MgII | ||||||
| 1.240 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.250 | MgII | ||||||
| 1.260 | MgII | ||||||
| 1.270 | MgII | ||||||
| 1.280 | MgII | ||||||
| 1.280 | MgII | ||||||
| 1.290 | MgII | ||||||
| 1.290 | MgII | ||||||
| 1.300 | MgII | ||||||
| 1.300 | MgII | ||||||
| 1.320 | MgII | ||||||
| 1.320 | MgII | ||||||
| 1.330 | MgII | ||||||
| 1.340 | MgII | ||||||
| 1.340 | MgII | ||||||
| 1.350 | MgII | ||||||
| 1.350 | MgII | ||||||
| 1.350 | MgII | ||||||
| 1.360 | MgII | ||||||
| 1.360 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.370 | MgII | ||||||
| 1.380 | MgII | ||||||
| 1.380 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.390 | MgII | ||||||
| 1.400 | MgII | ||||||
| 1.400 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.420 | MgII | ||||||
| 1.430 | MgII | ||||||
| 1.430 | MgII | ||||||
| 1.430 | MgII | ||||||
| 1.440 | MgII | ||||||
| 1.450 | MgII | ||||||
| 1.450 | MgII | ||||||
| 1.450 | MgII | ||||||
| 1.460 | MgII | ||||||
| 1.460 | MgII | ||||||
| 1.460 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.470 | MgII | ||||||
| 1.480 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.490 | MgII | ||||||
| 1.500 | MgII | ||||||
| 1.500 | MgII | ||||||
| 1.530 | MgII | ||||||
| 1.530 | MgII | ||||||
| 1.530 | MgII | ||||||
| 1.540 | MgII | ||||||
| 1.550 | MgII | ||||||
| 1.550 | MgII | ||||||
| 1.560 | MgII | ||||||
| 1.560 | MgII | ||||||
| 1.570 | MgII | ||||||
| 1.580 | MgII | ||||||
| 1.590 | MgII | ||||||
| 1.590 | MgII | ||||||
| 1.600 | MgII | ||||||
| 1.600 | MgII | ||||||
| 1.610 | MgII | ||||||
| 1.610 | MgII | ||||||
| 1.620 | MgII | ||||||
| 1.620 | MgII | ||||||
| 1.630 | MgII | ||||||
| 1.640 | MgII | ||||||
| 1.650 | MgII | ||||||
| 1.650 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.680 | MgII | ||||||
| 1.690 | MgII | ||||||
| 1.690 | MgII | ||||||
| 1.690 | MgII | ||||||
| 1.700 | MgII | ||||||
| 1.700 | MgII | ||||||
| 1.710 | MgII | ||||||
| 1.720 | MgII | ||||||
| 1.730 | MgII | ||||||
| 1.730 | MgII | ||||||
| 1.750 | MgII | ||||||
| 1.760 | MgII | ||||||
| 1.800 | MgII | ||||||
| 1.800 | CIV | ||||||
| 1.830 | MgII | ||||||
| 1.830 | CIV | ||||||
| 1.830 | MgII | ||||||
| 1.830 | CIV | ||||||
| 1.830 | MgII | ||||||
| 1.830 | CIV | ||||||
| 1.850 | MgII | ||||||
| 1.850 | CIV | ||||||
| 1.860 | MgII | ||||||
| 1.870 | MgII | ||||||
| 1.870 | CIV | ||||||
| 1.880 | MgII | ||||||
| 1.890 | MgII | ||||||
| 1.890 | CIV | ||||||
| 1.900 | MgII | ||||||
| 1.900 | CIV | ||||||
| 1.910 | MgII | ||||||
| 1.920 | CIV | ||||||
| 1.920 | MgII | ||||||
| 1.930 | CIV | ||||||
| 1.970 | CIV | ||||||
| 1.990 | CIV | ||||||
| 2.030 | CIV | ||||||
| 2.060 | CIV | ||||||
| 2.070 | CIV | ||||||
| 2.080 | CIV | ||||||
| 2.100 | CIV | ||||||
| 2.110 | CIV | ||||||
| 2.200 | CIV | ||||||
| 2.220 | CIV | ||||||
| 2.230 | CIV | ||||||
| 2.240 | CIV | ||||||
| 2.250 | CIV | ||||||
| 2.250 | CIV | ||||||
| 2.260 | CIV | ||||||
| 2.290 | CIV | ||||||
| 2.390 | CIV | ||||||
| 2.440 | CIV | ||||||